I am having trouble solving this problem because of all the negatives. Could someone help me with what all the steps are?
2(x-6)=5-{2x-[4(x-3)-5]}
Thanks!

Answer by Earlsdon(6294)

2(x-6) = 5-{2x-[4(x-3)-5]}
2x-12 = 5-{2x-[4x-12-5]}
2x-12 = 5-{2x-[4x-17]}
2x-12 = 5-{2x-4x+17}
2x-12 = 5-{-2x+17}
2x-12 = 5+2x-17
2x-12 = 2x-12
The solution is called "identity" and it means that all values of x will satisfy the equation.